21Pilots:

Which of the following is equivalent to 60^1/2 60/2 √60 1/60^2 1/√60

Tranquility:

There are few things you need to know before you can attempt the problem. First of all, \(\Huge a^{\left(\frac{b}{c}\right)}=\sqrt[c]{a^b}\) Also it's good to know that \(\Large\sqrt x \) is the same as \(\Large\sqrt[2]{x}\) With all of that known, what should the answer be?